OpenStudy (anonymous):

Sarah wants to buy the car that has a better mileage rating. Cars of Company A travel 20 miles per gallon, and cars of Company B travel 14 km per liter. [1 mile = 1.6 km; 1 gallon = 3.79 liters] Part A: Which company's car should Sarah buy? Justify your answer by comparing the mileages after converting to the same units. (5 points) Part B: If 10 gallons of gas is filled in the car which gives better mileage and you went on a drive and checked on the gas used every 30 miles, describe appropriate scales on the x-axis and y-axis to graph the rate at which the car uses gas. (5 points)

OpenStudy (shinalcantara):

all you have to do is convert the units say you convert 20miles per gallon to km/li \[20\frac{ mi }{ gal }(\frac{ 1.6km }{ 1 mi })(\frac{ 1gal }{ 3.79 li })\] try to compare it to 14km/li

OpenStudy (anonymous):

what i need help in the 3 a and b

OpenStudy (shinalcantara):

you need to compare their respective values in same unit of measurement for you to answer A.

OpenStudy (anonymous):

can u help me